You have $600 in an account which pays 4.4% compounded annually. How many additional dollars of interest would you earn over 5 years if you moved the money to an account earning 6.6​%?

How many additional dollars of interest would you earn over 5 years from the account that pays 6.6%?
Answer #1

Present value = Future value/(1+i)^n

i = interest rate per period

n= number of periods

Future value at 4.4% = 600 * (1+4.4%)^5

= 744.14

Future value at 6.6% = 600 * (1+6.6%)^5

= 825.91

additional dollars of interest = 825.91 - 744.14

= 81.77